# LeanVec: Search your vectors faster by making them fit

### v1

@@ Abstract
Modern deep learning models have the ability to generate high-dimensional vectors whose similarity reflects semantic resemblance. Thus, similarity search, i.e., the operation of retrieving those vectors in a large collection that are similar to a given query, has become a critical component of a wide range of applications that demand highly accurate and timely answers. In this setting, the high vector dimensionality puts similarity search systems under compute and memory pressure, leading to subpar performance. Additionally, cross-modal retrieval tasks have become increasingly common, e.g., where a user inputs a text query to find the most relevant images for that query. However, these queries often have different distributions than the database embeddings, making it challenging to achieve high accuracy. In this work, we present LeanVec, a framework that combines linear dimensionality reduction with vector quantization to accelerate similarity search on high-dimensional vectors while maintaining accuracy. We present LeanVec variants for in-distribution (ID) and out-of-distribution (OOD) queries. LeanVec-ID yields accuracies on par with those from recently introduced deep learning alternatives whose computational overhead precludes their usage in practice. LeanVec-OOD uses a novel technique for dimensionality reduction that considers the query and database distributions to simultaneously boost the accuracy and the performance of the framework even further (even presenting competitive results when the query and database distributions match). All in all, our extensive and varied experimental results show that LeanVec produces state-of-the-art results, with up to 3.7x improvement in search throughput and up to 4.9x faster index build time over the state of the art.
@@ 1 INTRODUCTION
Aguerrebere et al. (2023) have recently shown that, when properly implemented, graph search is bottlenecked by the memory bandwidth of the system, which is mainly consumed by fetching database vectors from memory in a random-like access pattern. Mainstream state-of-the-art vector quantization techniques (Andre et al., 2021; Guo et al., 2020) are incompatible with the random pattern. To reduce these access times in this setting, Aguerrebere et al. (2023) introduce a lightweight method, Locally-adaptive Vector Quantization (LVQ), that greatly accelerates the search and leads to state-of-the-art performance. Although LVQ removes the memory bottleneck in vectors of moderate dimensionality ($D\approx 128$), we observe increased memory bandwidth and computational pressure for higher dimensional (e.g., $D=512,768$) deep learning embedding vectors. Higher memory utilization drastically increases the memory latency (Srinivasan et al., 2009) to access each vector and results in suboptimal search performance. Even masterful placement of prefetching instructions in the software cannot hide the increased latency. These difficulties extend to the time-consuming procedure of constructing a graph-based index as construction speed is proportional to search speed. In a world where both academic and industrial deployments of similarity search overwhelmingly use deep learning embedding vectors, it is paramount to address this performance gap.
@@
In this work, we introduce LeanVec, a framework that combines linear dimensionality reduction with Locally-adaptive Vector Quantization (LVQ) to accelerate similarity search for high-dimensional vectors, making it suitable for applications with deep learning embedding vectors. We present LeanVec variants for the two main cases: in-distribution (ID) and out-of-distribution (OOD) queries. LeanVec’s compression lowers the required memory bandwidth and provides a proportional increase in search throughput. On the rqa-768-1M dataset (refer to LABEL:table:datasets for details), LeanVec compresses vectors by 9.6x (4.8x from the dimensionality reduction and 2x from LVQ) over the baseline of FP16 vectors and provides a 8.5x performance gain while consuming much less memory bandwidth (see Figure 1a). The performance gain increases even further as the dimensionality increases. For instance, in the standard gist-960-1M dataset, LeanVec shows ~12x improvement over the uncompressed FP16 vectors (see Figure 11 in the appendix).
We present the following contributions:
@@
  • For the ID case, LeanVec-ID, which combines principal component analysis (PCA) with LVQ, brings speedups of up to 3.6x for search. LeanVec-ID does not make any assumptions about the statistical distribution of the query vectors.
  • For the OOD case, we present a new linear dimensionality reduction technique, LeanVec-OOD, that finds the optimal projection subspaces for the dataset and a representative query set to reduce the errors in the similarity computations. The optimization in LeanVec-OOD is lightweight and fast, and we present a detailed analysis of its convergence. We show that LeanVec-OOD performs as good as LeanVec-ID in the ID case and is vastly superior in the OOD case.
@@ 2 LeanVec: a framework to accelerate similarity search for high-dimensional vectors
Notation. In this work, we denote vectors and matrices by lowercase and uppercase bold letters, respectively, e.g., ${\bm{\mathbf{v}}}\in{\mathbb{R}}^{n}$ and ${\bm{\mathbf{A}}}\in{\mathbb{R}}^{m\times n}$.
@@
where ${\bm{\mathbf{A}}},{\bm{\mathbf{B}}}\in{\mathbb{R}}^{d\times D}$ are orthonormal projection matrices, $d<D$, and $\operatorname{quant({\bm{\mathbf{v}}})}$ is a method to quantize each dimension in ${\bm{\mathbf{v}}}$. The projection matrices reduce the number of entries of the database vectors and the quantization reduces the number of bits per entry. The reduced memory footprint decreases the time it takes to fetch each vector from memory. Furthermore, the lower dimensionality alleviates the algorithm’s computational effort (i.e., requiring fewer fused multiply-add operations)
@@
The LeanVec framework is schematically depicted in Figure 1b and its constituents are described next. The computation of the projection matrices will be presented in sections 2.1 and 2.2. In the following, we refer to the vector set $\{\operatorname{quant}({\bm{\mathbf{B}}}{\bm{\mathbf{x}}}_{i})\,|\,{\bm{\mathbf{x}}}_{i}\in{\mathcal{X}}\}$ as primary vectors and to the set of $\{\operatorname{quant}({\bm{\mathbf{x}}}_{i})\,|\,{\bm{\mathbf{x}}}_{i}\in{\mathcal{X}}\}$ as secondary vectors.
@@
Graph construction. Only the primary vectors are used for graph construction. The secondary vectors are not used at this stage. Aguerrebere et al. (2023) had already analyzed the robustness of the graph construction to quantization with LVQ. Notably, our experimental results show that the robustness extends to a dimensionality reduction as well.
It is important to note that searches are an essential part of the graph construction process (Malkov & Yashunin, 2018; Fu et al., 2019). As such, our achieved search acceleration directly translates into graph construction acceleration, as shown in our experimental results.
LeanVec does not use dimensionality reduction to decrease the memory footprint of the similarity search index, but to accelerate its performance. LeanVec, in its current form, effectively increases the total footprint by keeping both primary and secondary vectors in memory. Without loss of generality and in pursuit of a reduced footprint, we could only store to $D-d$ dimensions for the secondary vectors (see the discussion in Section 2.1, which would remove the current overhead.
@@ 2.1 Dimensionality reduction for in-distribution similarity search
Let us begin with a few standard definitions. The Stiefel manifold is the set of orthonormal matrices, formally defined as
$\operatorname{St}(D,d)=\{{\bm{\mathbf{U}}}\in{\mathbb{R}}^{d\times D}\,|\,{\bm{\mathbf{U}}}{{\bm{\mathbf{U}}}}^{\top}={\bm{\mathbf{I}}}\}.$ (2)
Let $\left\|\bullet\right\|_{\text{op}}$ denote the standard operator norm, defined as $\left\|{\bm{\mathbf{A}}}\right\|_{\text{op}}=\sup\left\{\left\|{\bm{\mathbf{A}}}{\bm{\mathbf{v}}}\right\|_{2}/\left\|{\bm{\mathbf{v}}}\right\|_{2}\,|\,{\bm{\mathbf{v}}}\in{\mathbb{R}}^{D},{\bm{\mathbf{v}}}\neq{\bm{\mathbf{0}}}\right\}$.
The convex hull ${\mathcal{C}}$ of all orthonormal matrices in $\operatorname{St}(D,d)$ is the unit-norm ball of the operator norm, i.e.,
@@
In the in-distribution (ID) case, we compute the projection matrices from the set of database vectors ${\mathcal{X}}=\left\{{\bm{\mathbf{x}}}_{i}\in{\mathbb{R}}^{D}\right\}_{i=1}^{n}$. Let $d<D$. We use a matrix ${\bm{\mathbf{M}}}\in{\mathbb{R}}^{d\times D}$ to obtain a low-dimensional representation of the database vectors,
@@ 2.2 Query-aware dimensionality reduction for out-of-distribution similarity search
LeanVec-OOD allows suitable matrices for dimensionality reduction to be found and is specifically designed for the case where ${\mathcal{X}}$ and ${\mathcal{Q}}$ are not drawn from the same distribution. However, if ${\mathcal{X}}$ and ${\mathcal{Q}}$ are drawn from the same distribution, how does LeanVec compare to PCA? The following proposition addresses this question, showing that the LeanVec will perform similarly to PCA in the ID case (the proof is in Appendix A).
@@
Interestingly, for the searches performed as part of the graph construction process, database vectors are used as queries implying that, even for OOD use cases, the construction algorithm works with ID data. This highlights the importance of Proposition 1, ensuring that graph construction can be safely executed with LeanVec-OOD.
2.3 Optimizing the LeanVec-OOD objective function
@@
where
$f({\bm{\mathbf{A}}},{\bm{\mathbf{B}}})=\left\|{{\bm{\mathbf{Q}}}}^{\top}{{\bm{\mathbf{A}}}}^{\top}{\bm{\mathbf{B}}}{\bm{\mathbf{X}}}-{{\bm{\mathbf{Q}}}}^{\top}{\bm{\mathbf{X}}}\right\|_{F}^{2}.$ (10)
@@
The function $f$ in Equation 10 has continuous partial derivatives given by
$\tfrac{\partial}{\partial{{\bm{\mathbf{A}}}}}f({\bm{\mathbf{A}}},{\bm{\mathbf{B}}})$ $=2{\bm{\mathbf{B}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}{{\bm{\mathbf{B}}}}^{\top}{\bm{\mathbf{A}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}-2{\bm{\mathbf{B}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}},$ (13)
$\tfrac{\partial}{\partial{{\bm{\mathbf{B}}}}}f({\bm{\mathbf{A}}},{\bm{\mathbf{B}}})$ $=2{\bm{\mathbf{A}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}{{\bm{\mathbf{A}}}}^{\top}{\bm{\mathbf{B}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}-2{\bm{\mathbf{A}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}},$ (14)
where ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}={\bm{\mathbf{Q}}}{{\bm{\mathbf{Q}}}}^{\top}\in{\mathbb{R}}^{D\times D}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}={\bm{\mathbf{X}}}{{\bm{\mathbf{X}}}}^{\top}\in{\mathbb{R}}^{D\times D}$.
[…]
Equipped with these tools, we can pose the complete optimization procedure in LABEL:algo:leanvec. There, we update ${\bm{\mathbf{A}}}$ (resp. ${\bm{\mathbf{B}}}$) given a fixed ${\bm{\mathbf{B}}}$ (resp. ${\bm{\mathbf{A}}}$) by running one Frank-Wolfe update. The factor $\alpha\in(0,1)$, proposed by Wai et al. (2017) for the step size $\gamma=1/(t+1)^{\alpha}$, can be replaced by a line search to speed up the optimization. In our experiments we did not observe a need for such a performance tuning.
1 algorithm]algo:leanvec Let ${\bm{\mathbf{A}}}^{(0)},{\bm{\mathbf{B}}}^{(0)}\in{\mathcal{C}}$, e.g., ${\bm{\mathbf{A}}}^{(0)}\leftarrow{\bm{\mathbf{0}}}$ and ${\bm{\mathbf{B}}}^{(0)}\leftarrow{\bm{\mathbf{0}}}$;

### v2

# LeanVec: Searching vectors faster by making them fit

@@ Abstract
Modern deep learning models have the ability to generate high-dimensional vectors whose similarity reflects semantic resemblance. Thus, similarity search, i.e., the operation of retrieving those vectors in a large collection that are similar to a given query, has become a critical component of a wide range of applications that demand highly accurate and timely answers. In this setting, the high vector dimensionality puts similarity search systems under compute and memory pressure, leading to subpar performance. Additionally, cross-modal retrieval tasks have become increasingly common, e.g., where a user inputs a text query to find the most relevant images for that query. However, these queries often have different distributions than the database embeddings, making it challenging to achieve high accuracy. In this work, we present LeanVec, a framework that combines linear dimensionality reduction with vector quantization to accelerate similarity search on high-dimensional vectors while maintaining accuracy. We present LeanVec variants for in-distribution (ID) and out-of-distribution (OOD) queries. LeanVec-ID yields accuracies on par with those from recently introduced deep learning alternatives whose computational overhead precludes their usage in practice. LeanVec-OOD uses two novel techniques for dimensionality reduction that consider the query and database distributions to simultaneously boost the accuracy and the performance of the framework even further (even presenting competitive results when the query and database distributions match). All in all, our extensive and varied experimental results show that LeanVec produces state-of-the-art results, with up to 3.7x improvement in search throughput and up to 4.9x faster index build time over the state of the art.
@@ 1 INTRODUCTION
Aguerrebere et al. (2023) have recently shown that, when properly implemented, graph search is bottlenecked by the memory bandwidth of the system, which is mainly consumed by fetching database vectors from memory in a random access pattern (due to the complexity in globally ordering high-dimensional vectors and to the way any graph traversal works, i.e., hoping from one node to the other). Mainstream state-of-the-art vector quantization techniques are either specifically designed for a sequential access pattern (André et al., 2021; Guo et al., 2020), rendering them not suitable for graph search, or incur more expensive similarity calculations by increasing the number of memory accesses (Jégou et al., 2011) (more details in Section 4). To reduce these access times, Aguerrebere et al. (2023) introduce a lightweight method, Locally-adaptive Vector Quantization (LVQ), that greatly accelerates the search and leads to state-of-the-art performance. Although LVQ removes the memory bottleneck in vectors of moderate dimensionality ($D\approx 128$), we observe increased memory bandwidth and computational pressure for higher dimensional (e.g., $D=512,768$) deep learning embedding vectors. Higher memory utilization drastically increases the memory latency (Srinivasan et al., 2009) to access each vector and results in suboptimal search performance. Even masterful placement of prefetching instructions in the software cannot hide the increased latency. These difficulties extend to the time-consuming procedure of constructing a graph-based index as construction speed is proportional to search speed. In a world where both academic and industrial deployments of similarity search overwhelmingly use deep learning embedding vectors, it is paramount to address this performance gap.
@@
In this work, we introduce LeanVec, a framework that combines linear dimensionality reduction with Locally-adaptive Vector Quantization (LVQ) to accelerate similarity search for high-dimensional vectors, making it suitable for applications with deep learning embedding vectors. LeanVec is inscribed in the standard search-and-rerank paradigm popular in similarity search. We present LeanVec variants for the two main cases: in-distribution (ID) and out-of-distribution (OOD) queries. LeanVec’s compression lowers the required memory bandwidth and provides a proportional increase in search throughput. On the rqa-768-1M dataset (refer to LABEL:table:datasets for details), LeanVec compresses vectors by 9.6x (4.8x from the dimensionality reduction and 2x from LVQ) over the baseline of FP16 vectors and provides a 8.5x performance gain while consuming much less memory bandwidth (see Figure 1a). The performance gain increases even further as the dimensionality increases. For instance, in the standard gist-960-1M dataset, LeanVec shows ~12x improvement over the uncompressed FP16 vectors (see Figure 12 in the appendix). We present the following contributions:
@@
  • For the ID case, LeanVec-ID improves upon previous work using principal component analysis (PCA) (Jegou et al., 2010; Gong et al., 2012; Babenko & Lempitsky, 2014b; Wei et al., 2014) by combining it with LVQ, bringing search speedups of up to 3.6x over the state of the art. LeanVec-ID does not make any assumptions about the statistical distribution of the query vectors.
  • For the OOD case, we present a new linear dimensionality reduction technique, LeanVec-OOD, that finds the optimal projection subspaces for the dataset and a representative query set to reduce the errors in the similarity computations. We present two lightweight and fast optimization algorithms for LeanVec-OOD. We present a detailed convergence analysis for one of these variants (based on a non-convex Frank-Wolfe algorithm). We show that LeanVec-OOD performs as good as LeanVec-ID in the ID case and is vastly superior in the OOD case.
@@ 2 LeanVec: a framework to accelerate similarity search for high-dimensional vectors
Notation. We denote vectors/matrices by lowercase/uppercase bold letters, e.g., ${\bm{\mathbf{v}}}\in{\mathbb{R}}^{n}$ and ${\bm{\mathbf{A}}}\in{\mathbb{R}}^{m\times n}$.
@@
where ${\bm{\mathbf{A}}},{\bm{\mathbf{B}}}\in{\mathbb{R}}^{d\times D}$ are orthonormal projection matrices, $d<D$, and $\operatorname{quant({\bm{\mathbf{v}}})}$ is a method to quantize each dimension in ${\bm{\mathbf{v}}}$. The projection matrices reduce the number of entries of the database vectors and the quantization reduces the number of bits per entry. The reduced memory footprint decreases the time it takes to fetch each vector from memory. Furthermore, the lower dimensionality alleviates the algorithm’s computational effort (i.e., requiring fewer fused multiply-add operations). This approximation enables efficient inner product calculations with individual database vectors (no batch-processing required), which makes it ideal for the random memory-access pattern encountered in graph search.
@@
The LeanVec framework is schematically depicted in Figure 1b and its constituents are described next. The computation of the projection matrices will be presented in sections 2.1, 2.2, 2.3 and 2.4. In the following, we refer to the set $\{\operatorname{quant}({\bm{\mathbf{B}}}{\bm{\mathbf{x}}}_{i})\,|\,{\bm{\mathbf{x}}}_{i}\in{\mathcal{X}}\}$ as primary vectors and to the set $\{\operatorname{quant}({\bm{\mathbf{x}}}_{i})\,|\,{\bm{\mathbf{x}}}_{i}\in{\mathcal{X}}\}$ as secondary vectors.
@@
Graph construction. Only the primary vectors are used for graph construction. The secondary vectors are not used at this stage. Aguerrebere et al. (2023) had already analyzed the robustness of the graph construction to quantization with LVQ. Notably, our experimental results show that the robustness extends to a dimensionality reduction as well. It is important to note that searches are an essential part of the graph construction process (Malkov & Yashunin, 2018; Fu et al., 2019). As such, our achieved search acceleration directly translates into graph construction acceleration, as shown in our experimental results. See Appendix A for a discussion on graph construction and its acceleration.
LeanVec does not use dimensionality reduction to decrease the memory footprint of the similarity search index, but to accelerate its performance. LeanVec, in its current form, effectively increases the total footprint by keeping both primary and secondary vectors in memory. Without loss of generality and in pursuit of a reduced footprint, we could only store to $D-d$ dimensions for the secondary vectors (see the discussion in Section 2.1) which would remove the current overhead. Alternatively, other encodings (e.g., Douze et al., 2018) can be used for the secondary vectors.
@@ 2.1 Dimensionality reduction for in-distribution similarity search
Let us begin with a few standard definitions. The Stiefel manifold is the set of row-orthonormal matrices, formally defined as $\operatorname{St}(D,d)=\{{\bm{\mathbf{U}}}\in{\mathbb{R}}^{d\times D}\,|\,{\bm{\mathbf{U}}}{{\bm{\mathbf{U}}}}^{\top}={\bm{\mathbf{I}}}\}$. Let $\left\|\bullet\right\|_{\text{op}}$ denote the standard spectral norm, defined as $\left\|{\bm{\mathbf{A}}}\right\|_{\text{op}}=\sup\left\{\left\|{\bm{\mathbf{A}}}{\bm{\mathbf{v}}}\right\|_{2}/\left\|{\bm{\mathbf{v}}}\right\|_{2}\,|\,{\bm{\mathbf{v}}}\in{\mathbb{R}}^{D},{\bm{\mathbf{v}}}\neq{\bm{\mathbf{0}}}\right\}$. The convex hull ${\mathcal{C}}$ of all row-orthonormal matrices in $\operatorname{St}(D,d)$ is the unit-norm ball of the spectral norm, i.e.,
@@
In the in-distribution (ID) case, we compute the projection matrices from the set of database vectors ${\mathcal{X}}=\left\{{\bm{\mathbf{x}}}_{i}\in{\mathbb{R}}^{D}\right\}_{i=1}^{n}$. Let $d<D$. We use a matrix ${\bm{\mathbf{M}}}\in{\mathbb{R}}^{d\times D}$ to obtain the low-dimensional representation
@@ 2.2 Query-aware dimensionality reduction for out-of-distribution similarity search
LeanVec-OOD allows suitable matrices for dimensionality reduction to be found and is specifically designed for the case where ${\mathcal{X}}$ and ${\mathcal{Q}}$ are not drawn from the same distribution. However, if ${\mathcal{X}}$ and ${\mathcal{Q}}$ are drawn from the same distribution, how does LeanVec compare to PCA? The following proposition addresses this question, showing that the LeanVec will perform similarly to PCA in the ID case (the proof is in Appendix B).
@@
Interestingly, for the searches performed as part of the graph construction process, database vectors are used as queries implying that, even for OOD use cases, the construction algorithm works with ID data. Proposition 1 ensures that graph construction can be executed with LeanVec-OOD.
Efficiency. Developing the squared Frobenius norm, we can equivalently write Equation 7 as
$\min_{{\bm{\mathbf{A}}},{\bm{\mathbf{B}}}\in{\mathcal{C}}}\operatorname{Tr}\!\left({\bm{\mathbf{A}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}{{\bm{\mathbf{A}}}}^{\top}{\bm{\mathbf{B}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}{{\bm{\mathbf{B}}}}^{\top}+{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}-2{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}{{\bm{\mathbf{A}}}}^{\top}{\bm{\mathbf{B}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}\right),\quad\text{where}\quad{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}={\bm{\mathbf{Q}}}{{\bm{\mathbf{Q}}}}^{\top},{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}={\bm{\mathbf{X}}}{{\bm{\mathbf{X}}}}^{\top}$ (8)
Before solving this problem, we can precompute the $D\times D$ matrices ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$. This removes the optimization’s dependency in the number of database and query vectors and enables dealing with large data and query sets with great efficiency. Additionally, relying on the second-order statistics ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$ prevents from overfitting the query training set. Moreover, the error between a sample covariance matrix and its expectation converges very quickly (Koltchinskii & Lounici, 2017) with a growing sample size. We can thus safely use uniform subsampling to compute ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$, as observed in figures 15 and 16 of the appendix. We need a minimum of $D$ samples ($D$ query and $D$ database vectors) to ensure that these matrices are not artificially rank-deficient. Using $m=10^{4}$ queries and $n=10^{5}$ database vectors for training amounts to a 13x (20x) query oversampling for $D=512$ ($D=768$) over the minimum number of samples $D$.
2.3 Optimizing the LeanVec-OOD loss with a Frank-Wolfe algorithm
@@
$\text{where}\quad f({\bm{\mathbf{A}}},{\bm{\mathbf{B}}})=\left\|{{\bm{\mathbf{Q}}}}^{\top}{{\bm{\mathbf{A}}}}^{\top}{\bm{\mathbf{B}}}{\bm{\mathbf{X}}}-{{\bm{\mathbf{Q}}}}^{\top}{\bm{\mathbf{X}}}\right\|_{F}^{2}.$ (10)
@@
The function $f$ in Equation 10 has continuous partial derivatives given by (${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}},{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$ defined in Equation 8)
$\tfrac{\partial}{\partial{{\bm{\mathbf{A}}}}}f({\bm{\mathbf{A}}},{\bm{\mathbf{B}}})$ $=2{\bm{\mathbf{B}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}{{\bm{\mathbf{B}}}}^{\top}{\bm{\mathbf{A}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}-2{\bm{\mathbf{B}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}},\quad\text{and}\quad\tfrac{\partial}{\partial{{\bm{\mathbf{B}}}}}f({\bm{\mathbf{A}}},{\bm{\mathbf{B}}})$ $=2{\bm{\mathbf{A}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}{{\bm{\mathbf{A}}}}^{\top}{\bm{\mathbf{B}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}-2{\bm{\mathbf{A}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}.$ (13)
1 algorithm]algo:leanvec
2Let ${\bm{\mathbf{A}}}^{(0)},{\bm{\mathbf{B}}}^{(0)}\in{\mathcal{C}}$, e.g., ${\bm{\mathbf{A}}}^{(0)}\leftarrow{\bm{\mathbf{0}}}$ and ${\bm{\mathbf{B}}}^{(0)}\leftarrow{\bm{\mathbf{0}}}$;